OpenStudy (anonymous):

A line passes through the point (-3, -3) and has a slope of 1/2. What is the equation of the line? Is it y=1/2x - 3/2?

OpenStudy (anonymous):

You just have to use this formula\[y-\color{red}{y_o}=m*(x-\color{red}{x_o})\]Where\[(x_o,y_o)=(-3,-3)\]\[y-(-3)=\frac{1}{2}*(x-(-3))\]\[y+3=\frac{1}{2}x+\frac{3}{2}\]\[y=\frac{1}{2}x+\frac{3}{2}-3\]\[y=\frac{1}{2}x+\frac{3}{2}-\frac{6}{2}\]The answer is:\[\color{blue}{y=\frac{1}{2}x-\frac{3}{2}}\]